# Vacuum bubble and fissure formation in collective motion with competing attractive and repulsive forces Olivia Clifton1, Angel Chavez2, Antonio Madrigal3, Annie Warren2, Paige Yeung4, and Arnd Scheel2

### v1

@@ 1 Introduction
We study the continuum limit of large numbers of particles moving in the plane with positions $x_{j}\in\mathbb{R}^{2}$, $j=1,\ldots,N^{2}$ driven by an interaction potential $W_{{\mu}}$,
$\dot{x}_{j}=-\frac{1}{N^{2}}\sum_{j\neq\ell}^{N}\nabla W_{{\mu}}(x_{\ell}-x_{j}).$ (1.1)
@@
$u_{t}=\nabla\cdot\left(u\cdot(\nabla W_{{\mu}}*u)\right),\qquad x\in\mathbb{R}^{2};$ (1.2)
see for instance [11] and references therein. Letting the range of the repulsive potential converge to zero while increasing the strength appropriately, one can further approximate $W_{{\mu}}=\delta-{\mu}V$, which leads to the nonlinear diffusion equation with nonlocal interaction,
$u_{t}=\nabla\left(u\cdot\left(\nabla u-{\mu}\nabla V*u\right)\right),\qquad x\in\mathbb{R}^{2},$ (1.3)
@@
$u_{t}=\nabla\left(u\cdot\left(\nabla u-{\mu}\nabla V_{\mathrm{per}}*u\right)\right),\qquad x\in\Omega,$ (1.4)
with periodic boundary conditions. We focus on the
@@
$u_{t}=\varepsilon\Delta u+\nabla\left(u\cdot\left(\nabla u-{\mu}\nabla V_{\mathrm{per}}*u\right)\right),$ (1.7)
@@
We are interested in bifurcations from a constant distribution $u\equiv 1$. Clearly, this uniform density is always an equilibrium solution to (1.2) or (1.7), and one expects this state to be stable for weak attracting forces, $0<{\mu}\ll 1$. Increasing $\mu$}, one then indeed finds a destabilization and an ensuing phase transition, where densities will no longer be uniform but evolve into distributions with peaks, where particles cluster, and vacuum regions, where the density vanishes $u=0$. This transition curiously exhibits a vertical bifurcating branch, which leads to transitions that are reversible, that is, the bifurcation diagram does not display hysteresis, and to bifurcating patterns that exhibit vacuum regions at arbitrary small parameter distances from the bifurcation point; see for instance [5, 1, 3, 13]. In fact, we analyzed this transition in quite some detail in [13], finding universal expansions near the bifurcation point in quite general settings, as well as universal corrections for $0<\varepsilon\ll 1$. The work there focused on one-dimensional arrangements of particles, and motivates the study of the more complex two-dimensional arrangements that we are interested in here.
@@
In all of our results we focus on very simple potentials. In fact, due to periodization, potentials can be written as Fourier series over the fundamental domain, and we choose to keep only the first terms in these Fourier series, while retaining the lattice symmetry. In fact, starting with a radially symmetric potential $V$, the symmetrized potential $V_{\mathrm{per}}$ will have the lattice symmetries, that is, with $R,S$ from (1.6), $V_{\mathrm{per}}(R(x,y))=V_{\mathrm{per}}(x,y)$ and $V_{\mathrm{per}}(S(x,y))=V_{\mathrm{per}}(x,y)$, and of course $V_{\mathrm{per}}(x,y)=V_{\mathrm{per}}(x+e_{j,x},y+e_{j,y})$, $j=1,2$. To be specific, we consider the potentials
@@
Outline. We derive expansions for the size of vacuum bubbles and fissures near the bifurcation point in case of both square and hexagonal lattices in §2. The following section, §3, contains results on diffusive corrections, $\varepsilon\gtrsim 0$ in (1.7). In this setting, we also connect our findings to a more standard center-manifold analysis as for instance carried out in [4]. We study branches numerically and identify a secondary bifurcation mediating a change in topology and a transition between fissures and clusters in §4. We conclude with numerical results in the discrete setting, $N<\infty$, in §5 and a brief discussion, §6
@@ 2.1 The Vertical Branch
Solving for equilibria of (1.4), we find that $u\cdot\left(\nabla u-{\mu}\nabla V*u\right)=\nu$, a constant, which must in turn be 0. This is immediate when $u$ is not strictly positive. When $u>0$, we can divide by $u$ and find that the left-hand side is a gradient and non-zero, which is impossible for a periodic function. For a solution without vacuum, that is, for $u>0$, we then find
@@
${\mu}_{*}=\frac{1}{\pi^{2}},\quad u$ $=A_{0}+A_{1}\cos(x)\cos(y)+B_{1}\sin(x)\sin(y)\quad$ $\textrm{(sq.)},$ (2.2)
${\mu}_{*}=\frac{1}{\sqrt{3}\pi^{2}},\quad u$ $=A_{0}+A_{1}\cos\left(x+\frac{y}{\sqrt{3}}\right)+B_{1}\cos\left(x-\frac{y}{\sqrt{3}}\right)+C_{1}\cos\left(\frac{2y}{\sqrt{3}}\right).\quad$ (hex.) (2.3)
Fixing the average density to, say, 1, we may set $A_{0}=1$ and are left with two- and three-dimensional subspaces of solutions, respectively. These subspaces contain one-dimensional subspaces with larger symmetry (maximal isotropy) as follows. In the square symmetry case, $A_{1}=B_{1}$ corresponds to fissures, solutions depending on xyx-y, only, and $B_{1}=0$ corresponds to bubble solutions invariant under the 4-fold rotation $R$. Solutions are strictly positive, possessing no vacuum regions, as long as $\max(|A_{1}|,|B_{1}|)<A_{0}$.
In the hexagonal symmetry case, $A_{1}=B_{1}=C_{1}$ corresponds to bubble solutions with 6-fold rotational symmetry $R$, while fissures correspond to solutions where only one of $A_{1},B_{1},C_{1}$ is nonzero. Bubble solutions are strictly positive if $\frac{1}{3}A_{0}<A_{1},B_{1},C_{1}<\frac{2}{3}A_{0}$, fissure solutions for $|A_{1}|<1$, $B_{1}`=C_{1}=0$.
@@ 2.2 Vacuum Formation: Square Lattice Symmetry
We analyze solutions in the vicinity of the boundary of the vertical branch, when $\max(|A_{1}|,|B_{1}|)=A_{0}$. We look for non-negative solutions with support in $\Omega_{0}\subset\Omega$ that have average density $1$. They therefore satisfy
@@
Here, (2.4) follows from the evolution equation (1.4) by first concluding that $u\cdot\nabla(u-\mu V*u)=0$ and then, when $u\neq 0$, that $u-\mu V*u$ is constant. The continuity condition (2.6) follows from the fact that $u$ is a weak solution.
@@
Let $V(x,y)=2\cos(x)\cos(y)$, and set ${\mu}=\frac{1}{\pi^{2}}+{\widetilde{\mu}}$, with $\widetilde{\mu}$} sufficiently small. Then solutions $u$ to (2.4)-(2.6) with vacuum regions, that is, $\Omega_{0}\neq\Omega$, are of the form
@@
    $\ell=\frac{\pi}{2^{1/2}}\left(\frac{3}{2}{\widetilde{\mu}}\right)^{1/3}+\mathcal{O}({\widetilde{\mu}}^{\frac{2}{3}}),\qquad\mathcal{A}=4\pi^{2}\left(\frac{3}{2}{\widetilde{\mu}}\right)^{1/3}+\mathcal{O}({\widetilde{\mu}}^{\frac{2}{3}}).$
@@
    bubbles: We have $B_{1}=0$,
    $A_{0}=1-\frac{\pi^{2}}{4}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{\frac{3}{2}}),\qquad A_{1}=1+\frac{\pi^{\frac{3}{2}}}{\sqrt{2}}{\widetilde{\mu}}^{\frac{1}{2}}-\frac{\pi^{2}}{4}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{\frac{3}{2}}).$
@@
Figure 2.1: An illustration of vacuum bubbles (left) and fissures (right) in a square of width $2\pi$, comprising two (!) fundamental domains $\Omega$ of the square lattice. Blue regions represent areas of vacuum; $u$ is positive in the green regions $\Omega_{0}$. Note that there is one (complete) bubble and one fissure in each fundamental domain (boundaries of fundamental domain indicated by dashed lines); fissures have width $2\ell$ and total length $\sqrt{2}\pi$ in a fundamental domain, bubbles have radius $\ell$.
  • Proof . Throughout the following, we will use $\pi,\pi]^{2}$ as a domain of integration, noting that it contains exactly two fundamental domains $\Omega$.
    First, notice that (2.4) implies that for all $(x,y)$ where $u(x,y)>0$, we have
    $u(x,y)=\rho+{\mu}\iint_{\Omega_{0}}\cos(x-\xi)\cos(y-\eta)u(\xi,\eta)d\xi d\eta,$
    which, using trigonometric addition theorems gives the form (2.7). Therefore, all solutions with (or without) vacuum are found as solutions to a finite system of equations for $A_{0},A_{1}$, and $B_{1}$. We find solutions using a perturbative approach, starting with bubbles, $B_{1}=0$, $A_{1}\sim 0$, and $A_{0}\sim 1$. Substituting the ansatz $\left(A_{0}+A_{1}\cos(x)\cos(y)\right)_{+}$ into (2.4)-(2.6), we find
@@
    for a small parameter $z_{1}$. Our explicit ansatz then allows us to write the free-boundary condition (2.10) in polar coordinates about the centers $\{(0,\pm\pi),(\pm\pi,0)\}$ of the bubbles, as follows:
    $\begin{split}0&=A_{0}-A_{1}\left(1-\frac{r^{2}}{2}+\frac{r^{4}\cos^{2}(\theta)\sin^{2}(\theta)}{4}+\frac{r^{4}\cos^{4}(\theta)+r^{4}\sin^{4}(\theta)}{24}+\mathcal{O}(r^{6})\right)\\ &=\left(1+a_{0}\right)-\left(1+a_{0}+z_{1}^{2}\right)\left(1-\frac{r^{2}}{2}+r^{4}Q(\theta)+\mathcal{O}\left(r^{6}\right)\right),\end{split}$ (2.11)
    where $Q(\theta):=\frac{\cos^{2}(\theta)\sin^{2}(\theta)}{4}+\frac{\cos^{4}(\theta)+\sin^{4}(\theta)}{24}.$ The inner radius $\ell$ of the bubble can be identified at leading order as the value of $r$ which solves (2.11). We may now further scale $\ell=\ell_{1}z_{1}$, which yields, after some algebraic manipulation,
@@
    $0$ $=4\pi^{2}a_{0}-\iint_{\Omega^{c}}\left(1+a_{0}+\left(1+a_{0}+z_{1}^{2}\right)\cos(x)\cos(y)\right)dydx$
    $=4\pi^{2}a_{0}-4\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int_{0}^{z_{1}\ell_{1}(z_{1},a_{0}),\theta)}\left(-z_{1}^{2}+\frac{r^{2}}{2}+\mathcal{O}(a_{0}r^{2}+r^{4}+r^{2}z_{1}^{2})\right)r\ drd\theta$
    $=4\pi^{2}a_{0}-2\pi\left(-\frac{1}{2}z_{1}^{4}\right)+\mathcal{O}(a_{0}z_{1}^{4}+z_{1}^{6}).$
@@
    $0=$ $A_{1}-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(A_{1}\pi^{2})+\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}\cos(\xi)\cos(\eta)\left(A_{0}+A_{1}\cos(\xi)\cos(\eta)\right)d\eta d\xi$
@@
    $+4\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)\int_{0}^{\pi}\int_{0}^{z_{1}\ell_{1}(a_{0},z_{1},\theta)}\left(-1+\frac{r^{2}}{2}+\mathcal{O}(r^{4})\right)\left(A_{0}+A_{1}\left(-1+\frac{r^{2}}{2}+\mathcal{O}(r^{4})\right)\right)rdrd\theta$
    $=$ $-\pi^{2}{\widetilde{\mu}}\left(1+z_{1}^{2}-\frac{1}{2\pi}z_{1}^{4}\right)+\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)\left(2\pi z_{1}^{4}\right)+\mathcal{O}\left(z_{1}^{6}\right).$
    We thus see that $\widetilde{\mu}$} scales like $z_{1}^{4}$. We solve for $z_{1}$ as a function of ${\widetilde{\mu}}^{\frac{1}{4}}$ near $z_{1}={\widetilde{\mu}}^{\frac{1}{4}}=0$ by the implicit function theorem, obtaining
    $z_{1}=\left(\frac{\pi^{3}}{2}\right)^{\frac{1}{4}}{\widetilde{\mu}}^{\frac{1}{4}}+\mathcal{O}({\widetilde{\mu}}^{\frac{1}{2}}).$
@@
Remark 2.1 (Maximal Isotropy).
Note that the theorem demonstrates existence of branches with maximal isotropy, but does not explicitly exclude solutions with submaximal isotropy; see [10, 7] for background. Here, the isotropy subgroup is understood as a subgroup $G$ of the group $\Gamma$ of translations and rotations that map functions with lattice periodicty to functions with lattice periodicity, in this case a semi-direct product of translations in $x$ and $y$, $S^{1}\times S^{1}$, and the lattice isotropy $D_{4}$. In addition to basic lattice translations, the isotropy of bubbles is $D_{4}$, the isotropy of fissures is generated by $S^{1}$ and a reflection across fissures. Both are maximal subgroups of $\Gamma$. A rough calculation suggests that solutions with isotropy subgroups $G$ that are not maximal do not exist supercritically.

### v2

@@ 1 Introduction
We study the continuum limit of a large number of particles moving in the plane with positions $x_{j}\in\mathbb{R}^{2}$, $j=1,2,\ldots,M$ driven by an interaction potential $W_{{\mu}}$,
$\dot{x}_{j}=-\frac{1}{M}\sum_{1\leqslant j\leqslant M,j\neq\ell}\nabla W_{{\mu}}(x_{\ell}-x_{j}).$ (1.1)
@@
$u_{t}=\nabla\cdot\left(u\,\nabla(W_{{\mu}}*u)\right),\qquad x\in\mathbb{R}^{2},$ (1.2)
where convolution is defined through
$(W_{{\mu}}*u)(x)=\int_{y\in\mathbb{R}^{2}}W_{{\mu}}(x-y)u(y)dy;$
see for instance [3, 13] and references therein. Letting the range of the repulsive potential converge to zero while increasing the strength appropriately, one can further approximate $W_{{\mu}}=\delta-{\mu}V$, which leads to the nonlinear diffusion equation with nonlocal interaction,
$u_{t}=\nabla\cdot\left(u\,\nabla\left(u-{\mu}(V*u)\right)\right),\qquad u\in\mathbb{R},\ x\in\mathbb{R}^{2},$ (1.3)
@@
$u_{t}=\nabla\cdot\left(u\,\nabla\left(u-{\mu}(V_{\mathrm{per}}*u)\right)\right),\qquad x\in\Omega,$ (1.4)
with periodic boundary conditions, where now,
$(V_{\mathrm{per}}*u)(x)=\int_{y\in\Omega}V_{\mathrm{per}}(x-y)u(y)dy.$
Note the slight abuse of notation where convolution is denoted by the same symbol as in the whole plane. From now on, we shall always consider convolution in this periodic setting, thus removing any ambiguity. We focus on the
@@
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-{\mu}(V_{\mathrm{per}}*u)\right)\right),$ (1.7)
@@
We are interested in bifurcations from a constant distribution, which we choose as $u\equiv 1$ without loss of generality by scale-invariance of the equation. Clearly, this uniform density is always an equilibrium solution to (1.2) or (1.7), and one expects this state to be stable for weak attracting forces, $0<{\mu}\ll 1$. Increasing $\mu$}, one then indeed finds a destabilization and an ensuing phase transition, where densities will no longer be uniform but evolve into distributions with peaks, where particles cluster, and vacuum regions, where the density vanishes $u=0$. This transition curiously exhibits a vertical bifurcating branch, which leads to transitions that are reversible; that is, the bifurcation diagram does not display hysteresis; and to bifurcating patterns that exhibit vacuum regions at arbitrary small parameter distances from the bifurcation point; see for instance [7, 2, 5, 15]. We analyzed this transition in quite some detail in [15], finding universal expansions near the bifurcation point in quite general settings, as well as universal corrections for $0<\varepsilon\ll 1$. The work there focused on one-dimensional arrangements of particles, and motivates the study of the more complex two-dimensional arrangements that we are interested in here.
@@
In all of our results we focus on very simple potentials. In fact, due to periodization, potentials can be written as Fourier series over the fundamental domain, and we choose to keep only the first terms in these Fourier series, while retaining the lattice symmetry. Starting with a radially symmetric potential $V$, the symmetrized potential $V_{\mathrm{per}}$ will have the lattice symmetries; that is, with $R,S$ from (1.6), $V_{\mathrm{per}}(R(x,y))=V_{\mathrm{per}}(x,y)$ and $V_{\mathrm{per}}(S(x,y))=V_{\mathrm{per}}(x,y)$, and of course $V_{\mathrm{per}}(x,y)=V_{\mathrm{per}}(x+e_{j,x},y+e_{j,y})$, $j=1,2$. To be specific, we consider the potentials
@@
Outline. We derive expansions for the size of vacuum bubbles and fissures near the bifurcation point in case of both square and hexagonal lattices in §2. The following section, §3, contains results on diffusive corrections, $\varepsilon\gtrsim 0$ in (1.7). In this setting, we also connect our findings to a more standard center-manifold analysis as for instance carried out in [6], complemented there by global energetic considerations; see also [1] and [16] for generalizations to spherical geometry. We study branches numerically and identify a secondary bifurcation mediating a change in topology and a transition between fissures and clusters in §4. We conclude with numerical results in the discrete setting, $N<\infty$, in §5 and a brief discussion, §6.
@@ 2.1 The Vertical Branch
Solving for equilibria of (1.4), we find that $u\,\nabla\left(u-{\mu}(V*u)\right)=\nu\in\mathbb{R}^{2}$, a constant. If $u(x,y)=0$ at some point, then $\nu=0$. Otherwise, $\partial_{x}\left(u-{\mu}(V*u)\right)=\nu_{1}/u$ has a sign, which contradicts periodicity, unless $\nu_{1}=0$. With the same reasoning for the $y$-derivative, we conclude that $\nu=0$. For a solution without vacuum, that is, for $u>0$, we then find
@@
${\mu}_{*}=\frac{1}{\pi^{2}},\quad u=$ $A_{0}+A_{1}\cos(x)\cos(y)+B_{1}\sin(x)\sin(y)$
$+a_{1}\cos(x)\sin(y)+b_{1}\sin(x)\cos(y)\quad$ $\textrm{(sq.)},$ (2.2)
${\mu}_{*}=\frac{1}{\sqrt{3}\pi^{2}},\quad u=$ $A_{0}+A_{1}\cos\left(x+\frac{y}{\sqrt{3}}\right)+B_{1}\cos\left(x-\frac{y}{\sqrt{3}}\right)+C_{1}\cos\left(\frac{2y}{\sqrt{3}}\right)$
$+a_{1}\sin\left(x+\frac{y}{\sqrt{3}}\right)+b_{1}\sin\left(x-\frac{y}{\sqrt{3}}\right)+c_{1}\sin\left(\frac{2y}{\sqrt{3}}\right).\quad$ (hex.) (2.3)
We shall focus on solutions with $a_{1}=b_{1}=0$ in (2.2) and $a_{1}=b_{1}=c_{1}=0$ in (2.3) in the remainder of this paper. These subspaces consist of functions invariant under point reflections $(x,y)\mapsto(-x,-y)$. We only comment briefly on solutions outside of this fixed point subspace. Solutions with maximal isotropy all lie within these subspaces, possibly after a spatial translation; see Remarks 2.1 and 2.4, below.
Fixing the average density to, say, 1, we may set $A_{0}=1$ and are left with two- and three-dimensional subspaces of solutions, respectively. These subspaces contain one-dimensional subspaces with larger symmetry (maximal isotropy) as follows. In the square symmetry case, $A_{1}=B_{1}$ corresponds to fissures, solutions depending on xyx-y, only, and $B_{1}=0$ (or $A_{1}=0$) corresponds to bubble solutions invariant under the 4-fold rotation $R$. Solutions are strictly positive, possessing no vacuum regions, as long as $\max(|A_{1}|,|B_{1}|)<A_{0}$.
In the hexagonal symmetry case, $A_{1}=B_{1}=C_{1}$ corresponds to bubble solutions with 6-fold rotational symmetry $R$, while fissures correspond to solutions where only one of $A_{1},B_{1},C_{1}$ is nonzero. Bubble solutions are strictly positive if $\frac{1}{3}A_{0}<A_{1},B_{1},C_{1}<\frac{2}{3}A_{0}$, fissure solutions for $|A_{1}|<A_{0}=1$, $B_{1}=C_{1}=0$.
@@ 2.2 Vacuum Formation: Square Lattice Symmetry
We analyze solutions in the vicinity of the boundary of the vertical branch, when $\max(|A_{1}|,|B_{1}|)=A_{0}$. We look for non-negative solutions with support $\Omega_{0}\subset\Omega$ that have average density $1$. They therefore satisfy
@@
The continuity condition (2.6) follows from the fact that $u$ is a weak solution.
@@
Let $V(x,y)=2\cos(x)\cos(y)$, and set ${\mu}=\frac{1}{\pi^{2}}+{\widetilde{\mu}}$, with $\widetilde{\mu}$} sufficiently small. Then solutions $u$ to (2.4)-(2.6), invariant under point reflections u(x,y)=u(x,y)u(x,y)=u(-x,-y), with vacuum regions, that is, $\Omega_{0}\neq\Omega$, are of the form
@@
    $\ell=\frac{\pi}{2^{1/2}}\left(\frac{3}{2}{\widetilde{\mu}}\right)^{1/3}+\mathcal{O}({\widetilde{\mu}}^{2/3}),\qquad\mathcal{A}=4\pi^{2}\left(\frac{3}{2}{\widetilde{\mu}}\right)^{1/3}+\mathcal{O}({\widetilde{\mu}}^{2/3}).$
@@
    bubbles: $B_{1}=0$,
    $A_{0}=1-\frac{\pi^{2}}{4}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{3/2}),\qquad A_{1}=1+\frac{\pi^{3/2}}{\sqrt{2}}{\widetilde{\mu}}^{1/2}-\frac{\pi^{2}}{4}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{3/2}).$
@@
Figure 2.1: An illustration of vacuum bubbles (left) and fissures (right) in a square of width $2\pi$, comprising two (!) fundamental domains $\Omega$ of the square lattice. Blue regions represent areas of vacuum; $u$ is positive in the green regions $\Omega_{0}$. Note that there is one bubble and one fissure in each fundamental domain (boundaries of fundamental domain indicated by dashed lines); fissures have width $2\ell$ and total length $\sqrt{2}\pi$ in a fundamental domain, bubbles have radius $\ell$.
Proof . First, notice that (2.4) implies that for all $(x,y)$ where $u(x,y)>0$, we have
$u(x,y)=\rho+2{\mu}\iint_{\Omega_{0}}\cos(x-\xi)\cos(y-\eta)u(\xi,\eta)d\xi d\eta,$
which, using trigonometric addition theorems gives the form (2.7). Therefore, all solutions with (or without) vacuum are found as solutions to a finite system of equations for $A_{0},A_{1}$, and $B_{1}$. We find solutions using a perturbative approach, starting with bubbles, $B_{1}=0$, $A_{1}\sim 1$, and $A_{0}\sim 1$. Substituting the ansatz $\left(A_{0}+A_{1}\cos(x)\cos(y)\right)_{+}$ into (2.4)-(2.6), we find
@@
for a small parameter $z_{1}$. Our explicit ansatz then allows us to write the free-boundary condition (2.10) in polar coordinates about the center $(\pi,0)$ of the bubble, as follows:
$\begin{split}0&=A_{0}-A_{1}\left(1-\frac{r^{2}}{2}+\frac{r^{4}(3-\cos(4\theta))}{48}+\mathcal{O}(r^{6})\right)\\ &=\left(1+a_{0}\right)-\left(1+a_{0}+z_{1}^{2}\right)\left(1-\frac{r^{2}}{2}+r^{4}Q(\theta)+\mathcal{O}(r^{6})\right),\end{split}$ (2.11)
where $Q(\theta):=\frac{3-\cos(4\theta)}{48}.$The inner radius $\ell$ of the bubble can be identified at leading order as the value of $r$ which solves (2.11). We may now further scale $\ell=\ell_{1}z_{1}$, which yields, after some algebraic manipulation,
@@
$0$ $=2\pi^{2}-\iint_{\Omega}\left(A_{0}+A_{1}\cos(x)\cos(y)\right)dydx+\iint_{\Omega_{0}^{c}}\left(A_{0}+A_{1}\cos(x)\cos(y)\right)dydx$
$=2\pi^{2}-2\pi^{2}(1+a_{0})+\iint_{\Omega_{0}^{c}}\left(1+a_{0}+\left(1+a_{0}+z_{1}^{2}\right)\cos(x)\cos(y)\right)dydx$
$=-2\pi^{2}a_{0}+\int_{0}^{2\pi}\int_{0}^{z_{1}\ell_{1}(z_{1},a_{0}),\theta)}\left(-z_{1}^{2}+\frac{r^{2}}{2}+\mathcal{O}(a_{0}r^{2}+r^{4}+r^{2}z_{1}^{2})\right)r\ drd\theta$
$=-2\pi^{2}a_{0}+\pi\left(-\frac{1}{2}z_{1}^{4}\right)+\mathcal{O}(a_{0}z_{1}^{4}+z_{1}^{6}).$
@@
$0=$ $A_{1}-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(A_{1}\pi^{2})+\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}2\cos(\xi)\cos(\eta)\left(A_{0}+A_{1}\cos(\xi)\cos(\eta)\right)d\eta d\xi$
@@
$+2\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)\int_{0}^{2\pi}\int_{0}^{z_{1}\ell_{1}(a_{0},z_{1},\theta)}\left(-1+\frac{r^{2}}{2}+\mathcal{O}(r^{4})\right)\left(A_{0}+A_{1}\left(-1+\frac{r^{2}}{2}+\mathcal{O}(r^{4})\right)\right)rdrd\theta$
$=$ $-\pi^{2}{\widetilde{\mu}}\left(1+z_{1}^{2}-\frac{1}{2\pi}z_{1}^{4}\right)+\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)\left(2\pi z_{1}^{4}\right)+\mathcal{O}(z_{1}^{6}).$
We thus see that $\widetilde{\mu}$} scales like $z_{1}^{4}$. We solve for $z_{1}$ as a function of ${\widetilde{\mu}}^{1/4}$ near $z_{1}={\widetilde{\mu}}^{1/4}=0$ by the implicit function theorem, obtaining
$z_{1}=\left(\frac{\pi^{3}}{2}\right)^{1/4}{\widetilde{\mu}}^{1/4}+\mathcal{O}({\widetilde{\mu}}^{3/4}).$
@@
Remark 2.1 (Maximal Isotropy — square lattice).
Note that the theorem demonstrates existence of branches with maximal isotropy, but does not explicitly exclude solutions with submaximal isotropy; see [12, 9] for background. Here, the isotropy subgroup is understood as a subgroup $G$ of the group $\Gamma$ of translations and rotations that map functions with lattice periodicity to functions with lattice periodicity, in this case a semi-direct product of translations in $x$ and $y$, $S^{1}\times S^{1}$, and the lattice isotropy $D_{4}$. In addition to basic lattice translations, the isotropy of bubbles is $D_{4}$, the isotropy of fissures is generated by $S^{1}$ and a reflection across fissures. Both are maximal subgroups of $\Gamma$. A rough calculation suggests that solutions with isotropy subgroups $G$ that are not maximal do not exist supercritically. We also tried to find those numerically both in a finite-rank approximation that we analyze here and in the full PDE adding some artificial viscosity and determined that solutions with $D_{2}$ isotropy do not exist close to $\tilde{\mu}=0$. See also Remark 3.3 for diffusive corrections and §4.3 and §4.4 for $D_{2}$-symmetric solutions at finite, non-small $\tilde{\mu}$.